OpenStudy (anonymous):

how to solve 12x-2(4x-7)=42. Would like to see step process! Please! :)

OpenStudy (nincompoop):

disitribute -2 to 4x - 7 first

OpenStudy (anonymous):

so it would be 12x-8x+14=42 ?

OpenStudy (nincompoop):

yeah then simplify

OpenStudy (anonymous):

Thank you soo much! :) Would you happen to know how to figure out 4[2-(2x-3)]=2(8-3x)

OpenStudy (nincompoop):

yeah

OpenStudy (nincompoop):

first work whatever is in the innermost part

OpenStudy (nincompoop):

-(2x-3) is simply -1(2x-3) distribute or it's a matter of change in signs

OpenStudy (anonymous):

so 4[-4x+6]=16-6x ?

OpenStudy (nincompoop):

2-2x+3

OpenStudy (nincompoop):

then it becomes 4(2-2x+3) or 4(-2x-1)

OpenStudy (nincompoop):

you are not distributing the 2 on the left hand side of the equation, there is a minus sign after the 2 so it is a subtraction operation not multiplication

OpenStudy (anonymous):

okay.